# Vulnerability Webs:

### v2

@@
Systemic Risk in Software Networksthanks: The order of authors is alphabetical. fritzc@tcd.ie, co-pierre.georg@uct.ac.za, angelo.mele@jhu.edu, and mus47@psu.edu. This paper supersedes a previous version titled "A Strategic Model of Software Dependency Networks". We wish to thank Bryan Graham, Stephane Bonhomme, Louise Laage, Matt Jackson, Eric Auerbach for helpful suggestions. We also thank participants at 2024 Cowles Foundation Conference in Econometrics, ESIF 2024 Conference on ML+AI, Empirical Methods in Game Theory Workshop, MEG 2024, EC24, MGOW 2024, Network Science in Economics 2023, 2023 BSE Summer Forum Networks Workshop, SEA 2023, Econometrics of Peer Effects and Networks, and seminar participants at Georgetown University, University of Washington, Virginia Tech, Boise State University, and the University of Pennsylvania for helpful comments and suggestions. The authors acknowledge support from the German Research Foundation award DFG FR 4768/1-1 (CF), Ripple’s University Blockchain Research Initiative (CPG), the U.S. Department of Defense award ARO W911NF-21-1-0335 (MS, CF), and the U.S. National Science Foundation awards NSF DMS-1513644 and NSF DMS-1812119 (MS). Mele acknowledges support from an IDIES Seed Grant from the Institute for Data Intensive Engineering and Science at Johns Hopkins University and partial support from NSF grant SES 1951005.
@@
Modern software development is a collaborative effort that re-uses existing code to reduce development and maintenance costs. This practice exposes software to vulnerabilities in the form of undetected bugs in direct and indirect dependencies, as demonstrated by the Crowdstrike and HeartBleed bugs. The economic costs resulting from such vulnerabilities can be staggering. We study a directed network of 52,897 software dependencies across 16,102 Python repositories, guided by a strategic model of network formation that incorporates both observable and unobservable heterogeneity. Using a scalable variational approximation of the conditional distribution of unobserved heterogeneity, we show that outsourcing code to other software packages by creating dependencies generates negative externalities. Modeling the propagation of risk in networks of software packages as an epidemiological process, we show that increasing protection of dependencies based on popular heuristics is ineffective at reducing systemic risk. By contrast, AI-assisted coding enables developers to replace dependencies with in-house code and reduces systemic risk.
@@ 1 Introduction
On July 19, 2024 a faulty update to the software Crowdstrike—a security software adopted by many companies—created worldwide panic and disruptions, paralyzing domestic and international air traffic for hours, and disrupting operations at banks, hospitals, and hotel chains. [^1] See https://www.nytimes.com/2024/07/19/business/dealbook/tech-outage-crowdstrike-microsoft.html All told, the update affected 8.5 million Windows machines worldwide, each of which had to be manually rebooted to patch up the system and restart operations. The faulty update of the software Crowdstrike was not the first instance highlighting systemic risk in software systems, and it will not be the last: e.g., in 2014 the HeartBleed bug—a vulnerability of the OpenSSL authentication library used by many websites—compromised the data of millions of users of websites such as Eventbrite, OkCupid, in addition to disrupting operations at the FBI and hospitals. Such disruptions can generate staggering economic costs: Heartbleed, for example, resulted in at least \$currency-dollar\\$\$500 million in damages. According to the risk management company Parametrix, the disruption caused by Crowdstrike cost Fortune 500 companies \$currency-dollar\\$\$5.4 billion. [^2] See Parametrix (2024).
@@
The risk of these disruptions is deeply embedded in the structure of software systems. Modern software development is a collaborative effort that produces sophisticated software, making extensive use of already existing software packages and libraries (Schueller et al., 2022). This results in a complex network of dependencies among software packages, best described as dependency graphs. [^3] Dependency graphs have been studied for many different programming languages using information from package managers. For an empirical comparison of the most common ones, see Decan et al. (2019). While this re-use of code produces significant efficiency gains for software developers (Lewis et al., 1991; Barros-Justo et al., 2018), it also increases the risk that a vulnerability in one software package renders a large number of other packages equally vulnerable. In line with this view, the number of vulnerabilities listed in the Common Vulnerabilities and Exposure Database has increased almost twelve-fold between 2001 and 2019.444The Common Vulnerabilities and Exposure database is a public reference for known information-security vulnerabilities which feeds into the U.S. National Vulnerability Database. See https://cve.mitre.org/. The reported number of incidents has increased from 3,591 for the three years from 1999 to 2001 to 43,444 for the three-year period from 2017 to 2019. A recent estimate by Krasner (2018) put the cost of losses from software failures at over \$1 trillion, up from \$59 billion in 2002 estimated by the National Institute of Standards and Technology (2002). [^5] Industry group Cybersecurity Ventures estimates that the damages incurred by all forms of cyber crime, including the cost of recovery and remediation, totalled USD 6 Trillion in 2021, and could reach USD 10.5 Trillion annually by 2025 (Source).
@@
Because vulnerabilities can potentially spread from one software package to all packages depending on it, coders can create an externality when deciding to re-use code from existing packages rather than implementing the required functionality themselves. Whether this is the case and how large the externality is, is ultimately an empirical question.
@@
We model coders’ decisions to create a dependency to another software package as an equilibrium game of network formation with observed and unobserved heterogeneity. Developers weigh costs and benefits of forming links and revise these links at random times, when there is a need for an update. Our equilibrium characterization provides the likelihood of observing a particular architecture of the software dependency network in the long run. Using data on open source software projects contained in Python’s software package manager Pypi, we obtain $52,897$ dependencies among $16,102$ repositories. We find evidence that a maintainer’s decision to allow a dependency exerts a negative externality on other maintainers. In principle, the presence of this externality discourages the formation of dependencies and partially contributes to the sparsity of the network. Our result implies that it is particularly important to ensure that such packages are free of vulnerabilities that potentially can affect a large number of other software packages.
@@
We can use our model to study the effect of interventions that reduce vulnerabilities of libraries that are likely to affect a large subset of the software ecosystem. We assume that vulnerabilities, coding errors or bugs spread according to an epidemic susceptible-infected-recovered (SIR) model. Our simulations show that strategies to decrease systemic risk of contagion based on interventions that target nodes according to measures of centrality or expected risk of spread, are not very effective in mitigating the average systemic risk. On the other hand we observe that the introduction of AI-assisted coding may decrease the incentives to form dependencies through a decrease in costs of producing code. This process of decreasing the density of the network will mechanically decrease the average systemic risk. We show that even moderate productivity gains from AI translate into a significant flattening of the risk profile for the software network.
@@
Open source software is an ideal laboratory to study software dependencies. Since the actual source code of a software package can be inspected and modified by others, code reuse is not only possible, it is encouraged. Modern software packages are not developed in isolation, but within an ecosystem of interdependent packages. A package manager keeps track of code dependencies and ensures that these are automatically resolved so that when a user installs one package, all of the package’s dependencies are also installed and automatically satisfied. There are many programming languages that provide software using package managers. We restrict our analysis to repositories of software packages written in the popular programming language Python and managed by the Pypi package manager.666 PyPI stands for Python Package Index and is Python’s official third-party software repository. See here for the official documentation. Python is a modern programming language that is very popular in the software development community. We use data from libraries.io, which, for each software repository includes a full list of all dependencies.
@@
We model the development of software as a process where coders reuse existing code, creating a network of dependencies among software packages—organized in repositories, which can include several closely related software packages—in the process. This is typical for today’s prevalent object-oriented software development paradigm. Our model describes a system of $N$ individual software repositories, characterized by observable and unobservable types. Each repository is managed by a group of coders who decide which dependencies to form. Coders obtain a net benefit of linking to another package which depends on how active this dependency is maintained, how mature, popular, and large it is. We also allow a coders’ utility to be affected by a local, i.e. type-specific, externality: since dependent packages have dependencies themselves, they are susceptible to vulnerabilities imported from other packages. So we assume that coders care about the direct dependencies of the packages they link to. This specification excludes externalities that are more than two links away as in the network formation models of DePaula et al. (2018), Mele (2017) and Mele and Zhu (2023). The network of dependencies forms over time and in each period a randomly selected package needs an update, so the coders decide whether to form a link or update the software in-house. Before updating the link, the package receives a random match quality shock. We characterize the equilibrium distribution over networks as a mixture of exponential random graphs (Schweinberger and Handcock, 2015; Mele, 2022), which can be decomposed into within- and between-types contribution to the likelihood.
@@
Our main result is that coders exert a negative externality on other coders when creating a dependency on another package. Other coders are then also less likely to create a dependency with that package. As a result, the network is relatively sparse. The interdependence among libraries implies that there exists a risk that a vulnerability in a single package has large adverse consequences for the entire ecosystem. One example of how a vulnerability in one package impacted a significant part of critical web infrastructure is Heartbleed, the infamous bug in the widely used SSL/TLS cryptography library, resulting from an improper input variable validation (Durumeric et al., 2014). When Heartbleed was disclosed in 2014, it rendered up to 55% of secure webservers vulnerable to data theft, including of the Canadian Revenue Agency, and Community Health Systems, a large U.S. hospital chain. Our model provides a further argument for ensuring the security of highly interdependent software packages. Not only can a vulnerability in such a package affect a relatively large fraction of the entire ecosystem, because of the externality we identify, it is also likely that this fraction increases as time goes on.
@@
Another important question in the study of network formation processes is whether linked packages are similar or dissimilar with respect to their covariates. We find evidence of similarity among the same type of software packages in terms of their Size, Popularityand Maturityas well as between different-type software packages in terms of their Popularity. The effect of similarity decreases with the (log of) the number of packages of the same unobservable type. We further find evidence of dis-similarity between different-type software packages in terms of their Maturity  and Size. In other words, mature, popular and large software packages of one type are likely to depend on similar packages of the same type, but on less mature and smaller software packages of another type. This is intuitive, because mature, popular, and large packages are likely to have a large user base with high expectations of the software and coders try to satisfy this demand by providing sophisticated functionality with the help of mature, popular, and large software packages of the same type. These packages are often built using a lot of low-level functionality from large but fairly generic libraries, which is why large popular and mature software packages of one type depend on larger, but less popular and less mature software packages of another type. One example of this is the inclusion of a payment gateway in an e-commerce application. The e-commerce application itself can be sophisticated and complex, like ebay is in the provision of their auction mechanism. For such an application it is particularly valuable to provide users with the ability to use a variety of different payment methods, including credit card, paypal, or buy now pay later solutions like Klarna. Our result also aligns with the recent trend in software development away from large monolithic applications and towards interconnected microservices (Traore et al., 2022).
Lastly, we adapt the susceptible-infected-recovered (SIR) model from epidemiology to examine the spread of vulnerabilities in the Python Pypi dependency graph. We measure a repository’s $k$-step systemicness as the number of downstream packages that are potentially rendered vulnerable by a vulnerability or bug in the upstream repository.
@@
We then study the efficacy of targeted interventions to mitigate vulnerability spread, drawing parallels with vaccination strategies in public health. By focusing on securing the most critical nodes—determined according to in-degree, expected fatality or a combination of betweenness centrality and expected fatality—we assess the potential to significantly reduce the risk of vulnerability contagion.
@@
Our findings suggest that protecting even ten percent of the most critical repositories may reduce the average systemic risk, but cannot completely flatten out the risk profile of the ecosystem. On the other hand, we argue that the introduction of AI-assisted coding may have a larger effect on reducing systemic risk. Recent work has shown that AI tools such as GitHub Copilot can increase productivity of software developers (Peng et al., 2023; Edelman et al., 2023; Noy and Zhang, 2023). In our model, this corresponds to a decrease of the cost of producing software in-house, without relying on creating dependencies or creating fewer dependencies. This can be modeled as an increase of the relative cost of a dependency. By simulating different magnitudes of the productivity improvement from AI-assisted coding, we show that the average systemic risk profile is flattened out by an increase of cost of merely five percent. The mechanism through which this risk reduction is achieved is the brute-force decrease in density for the network implied by increases in costs of dependencies. Overall, our results help better understand the various driving forces and motifs in software development and how they shape the network of software dependencies.
@@
Our model is able to estimate the magnitude of externalities as well as homophily (Currarini et al., 2010; Jackson, 2008; Chandrasekhar, 2016), the tendency of individuals to form links to similar individuals. Furthermore, our model is able to detect heterophily (or competition) among maintainers. More specifically, we allow the homophily to vary by unobservables, while in most models the homophily is estimated only for observable characteristics (Currarini et al., 2010; Schweinberger and Handcock, 2015; DePaula et al., 2018; Chandrasekhar, 2016; Graham, 2020).

### v3

@@
Systemic Risk in Software Networksthanks: Contact: fritzc@tcd.ie, co.georg@fs.de, angelo.mele@jhu.edu, and mus47@psu.edu. We wish to thank Bryan Graham, Stephane Bonhomme, Louise Laage, Matt Jackson, Eric Auerbach for helpful suggestions. We also thank participants at 2024 Cowles Foundation Conference in Econometrics, ESIF 2024 Conference on ML+AI, Empirical Methods in Game Theory Workshop, MEG 2024, EC24, MGOW 2024, Network Science in Economics 2023, 2023 BSE Summer Forum Networks Workshop, SEA 2023, Econometrics of Peer Effects and Networks, and seminar participants at Georgetown University, University of Washington, Virginia Tech, Boise State University, and the University of Pennsylvania for helpful comments and suggestions. The authors acknowledge support from the German Research Foundation award DFG FR 4768/1-1 (CF), Ripple’s University Blockchain Research Initiative (CPG), the U.S. Department of Defense award ARO W911NF-21-1-0335 (MS, CF), and the U.S. National Science Foundation awards NSF DMS-1513644 and NSF DMS-1812119 (MS). Mele acknowledges support from an IDIES Seed Grant from the Institute for Data Intensive Engineering and Science at Johns Hopkins University and partial support from NSF grant SES 1951005.
@@
Software development relies on code reuse to minimize costs, creating vulnerability risks through dependencies with substantial economic impact, as seen in the Crowdstrike and HeartBleed incidents. We analyze 52,897 dependencies across 16,102 Python repositories using a strategic network formation model incorporating observable and unobservable heterogeneity. Through variational approximation of conditional distributions, we demonstrate that dependency creation generates negative externalities. Vulnerability propagation, modeled as a contagion process, shows that popular protection heuristics are ineffective. AI-assisted coding, on the other hand, offers an effective alternative by enabling dependency replacement with in-house code.
@@ 1 Introduction
On July 19, 2024 a faulty update to the software Crowdstrike—a security software adopted by many companies—created worldwide panic and disruptions, paralyzing domestic and international air traffic for hours, and disrupting operations at banks, hospitals, and hotel chains. [^1] See https://www.nytimes.com/2024/07/19/business/dealbook/tech-outage-crowdstrike-microsoft.html All told, the update affected 8.5 million Windows machines worldwide, each of which had to be manually rebooted to patch up the system and restart operations. The faulty update of the Crowdstrike software was not the first instance highlighting systemic risk in software systems, and it will not be the last: e.g., in 2014 the Heartbleed bug—a vulnerability of the OpenSSL authentication library used by many websites—compromised the data of millions of users of websites such as Eventbrite, OkCupid, in addition to disrupting operations at the FBI and hospitals. Such disruptions can generate staggering economic costs: Heartbleed, for example, resulted in at least \$currency-dollar\\$\$500 million in damages. According to the risk management company Parametrix, the disruption caused by Crowdstrike cost Fortune 500 companies \$currency-dollar\\$\$5.4 billion. [^2] See Parametrix (2024).
@@
The risk of these disruptions is deeply embedded in the structure of software systems. Modern software development is a collaborative effort that produces sophisticated software by making extensive use of already existing code (Schueller et al., 2022). This results in a complex network of dependencies among software packages, best described as dependency graphs. [^3] Dependency graphs have been studied for many different programming languages using information from package managers. For an empirical comparison of the most common ones, see Decan et al. (2019). While this re-use of code produces significant efficiency gains for software developers (Lewis et al., 1991; Barros-Justo et al., 2018), it also increases the risk that a vulnerability in one software package renders a large number of other packages equally vulnerable. In line with this view, the number of vulnerabilities listed in the Common Vulnerabilities and Exposure Database has increased almost twelve-fold between 2001 and 2019.444The Common Vulnerabilities and Exposure database is a public reference for known information-security vulnerabilities which feeds into the U.S. National Vulnerability Database. See https://cve.mitre.org/. The reported number of incidents has increased from 3,591 for the three years from 1999 to 2001 to 43,444 for the three-year period from 2017 to 2019. An estimate by Krasner (2018) put the cost of losses from software failures at over \$1 trillion, up from \$59 billion in 2002 estimated by the National Institute of Standards and Technology (2002). [^5] Industry group Cybersecurity Ventures estimates that the damages incurred by all forms of cyber crime, including the cost of recovery and remediation, totalled USD 6 Trillion in 2021, and could reach USD 10.5 Trillion annually by 2025 (Source).
@@
Because vulnerabilities can potentially spread from one software package to all packages depending on it, coders can create an externality when deciding to re-use code from existing packages rather than implementing the required functionality themselves. Whether this is the case and how large the externality is is ultimately an empirical question.
@@
We model coders’ decisions to create a dependency to another software package as an equilibrium game of network formation with observed and unobserved heterogeneity. [^6] Formally, we aggregate software packages that belong to the same code repository. Each package belongs to a single repository, but each repository can contain multiple packages. Repositories are typically managed by a small team of maintainers who decide which code is added to the repository. Developers weigh the costs and benefits of forming links and revise these links at random times when there is a need for an update. Our equilibrium characterization provides the likelihood of observing a particular architecture of the software dependency network in the long run. Using data on open source software projects contained in Python’s software package manager Pypi, we obtain $52,897$ dependencies among $16,102$ repositories. We find evidence that a maintainer’s decision to allow a dependency exerts a negative externality on other maintainers. In principle, the presence of this externality discourages the formation of dependencies and partially contributes to the sparsity of the network. Our result implies that it is particularly important to ensure that the code in these repositories is free from vulnerabilities that potentially affect the code in a large number of other repositories.
@@
We can use our model to study the effect of interventions that reduce code vulnerabilities that are likely to affect a large subset of the software ecosystem. We assume that vulnerabilities, coding errors, or bugs spread according to an epidemiological contagion model. Our simulations show that strategies to decrease the risk of contagion based on interventions targeting nodes according to measures of centrality or expected spread risk are not very effective in mitigating the average risk of contagion. On the other hand, we observe that the introduction of AI-assisted coding may decrease the incentives to form dependencies through a decrease in the costs of producing code. This process of decreasing the density of the network will mechanically decrease the average risk of contagion. We show that even moderate productivity gains from AI translate into a significant flattening of the risk profile for the software network.
@@
Open-source software is an ideal laboratory to study software dependencies. Since the actual source code of software in a repository can be inspected and modified by others, code reuse is not only possible, it is encouraged. Modern software is not developed in isolation, but within an ecosystem of interdependent packages organized in code repositories. A package manager keeps track of code dependencies and ensures that these are automatically resolved so that when a user installs one package from a repository, all of the package’s dependencies are also installed and automatically satisfied. There are many programming languages that provide software using package managers. We focus on repositories of software packages written in the popular Python programming language and managed by the Pypi package manager.777 PyPI stands for Python Package Index and is Python’s official third-party software repository. See here for the official documentation. Python is a modern programming language that is very popular in the software development community. We use data from libraries.io, which, for each software repository, includes a full list of all dependencies.
@@
We model the development of software as a process where maintainers reuse existing code, creating a network of dependencies among software packages—organized in repositories of several closely related software packages—in the process. This is typical for today’s prevalent object-oriented software development paradigm. Our model describes a system of $N$ individual software repositories, characterized by observable and unobservable types. Each repository is managed by a single maintainer who decides which code to add to the repository and, consequently, which dependencies to form. Maintainers obtain a net benefit of linking to another package which depends on how active this dependency is maintained, how mature, popular, and large it is. We also allow a maintainers’ utility to be affected by a local, i.e. type-specific, externality: Since dependent packages have dependencies themselves, they are susceptible to vulnerabilities imported from other packages. So we assume that maintainers care about the direct dependencies of the packages they link to. This specification excludes externalities that are more than two links away, as in the network formation models of DePaula et al. (2018), Mele (2017) and Mele and Zhu (2023). The network of dependencies forms over time and in each period a randomly selected package needs an update, so the maintainers decide whether to form a link or update the software in-house. Before updating the link, the package receives a random match quality shock. We characterize the equilibrium distribution over networks as a mixture of exponential random graphs (Schweinberger and Handcock, 2015; Mele, 2022), which can be decomposed into within- and between-types contribution to the likelihood.
@@
Our main result is that maintainers exert a negative externality on other maintainers when creating a dependency on another repository: Other maintainers are then less likely to create a dependency on that repository. As a result, the network is relatively sparse. The interdependence among libraries implies that there exists a risk that a vulnerability in a single package has large adverse consequences for the entire ecosystem. An example of how a vulnerability in one package affected a significant part of critical web infrastructure is Heartbleed, the infamous bug in the widely used SSL/TLS cryptography library, resulting from incorrect validation of the input variable (Durumeric et al., 2014). When Heartbleed was disclosed in 2014, it made up to 55% of secure webservers vulnerable to data theft, including the Canadian Revenue Agency and Community Health Systems, a large US hospital chain. Our model provides a further argument for ensuring the security of highly interdependent software packages. Not only can a vulnerability in such a package affect a relatively large fraction of the entire ecosystem, because of the externality we identify, it is also likely that this fraction increases as time goes on.
@@
Another important question in the study of network formation processes is whether linked packages are similar or dissimilar with respect to their covariates. We find evidence of similarity among the same type of software packages in terms of their Size, Popularity and Maturity as well as between different-type software packages in terms of their Popularity. The effect of similarity decreases with the (log of) the number of packages of the same unobservable type. We further find evidence of dis-similarity between different-type software packages in terms of their maturity and size. In other words, mature, popular, and large software packages of one type are likely to depend on similar packages of the same type, but on less mature and smaller software packages of another type. This is intuitive because mature, popular, and large packages are likely to have a large user base with high expectations of the software, and coders try to satisfy this demand by providing sophisticated functionality with the help of mature, popular, and large software packages of the same type. These packages are often built using a lot of low-level functionality from large but fairly generic libraries, which is why large popular and mature software packages of one type depend on larger, but less popular and less mature software packages of another type. One example of this is the inclusion of a payment gateway in an e-commerce application. The e-commerce application itself can be sophisticated and complex, like ebay is in the provision of their auction mechanism. For such an application, it is particularly valuable to provide users with the ability to use a variety of different payment methods, including credit card, paypal, or buy now pay later solutions like Klarna. Our result also aligns with the recent trend in software development away from large monolithic applications and towards interconnected microservices (Traore et al., 2022).
@@
Lastly, we adapt a simple epidemiological contagion model to examine the spread of vulnerabilities in the Pypi dependency graph. We measure a repository’s $k$-step systemicness as the number of downstream packages that are potentially rendered vulnerable by a vulnerability or bug in the upstream repository. We use this model to study the efficacy of targeted interventions to mitigate vulnerability spread, drawing parallels with vaccination strategies in public health. By focusing on securing the most critical nodes–determined according to in-degree, expected fatality, or a combination of betweenness centrality and expected fatality–we assess the potential to significantly reduce the risk of vulnerability contagion.
@@
Our findings suggest that protecting even ten percent of the most critical repositories may reduce the average systemic risk but cannot completely flatten out the risk profile of the ecosystem. On the other hand, we argue that the introduction of AI-assisted coding may have a greater effect on reducing systemic risk. Recent work has shown that AI tools such as GitHub Copilot can increase the productivity of software developers (Peng et al., 2023; Edelman et al., 2023; Noy and Zhang, 2023). In our model, this corresponds to a decrease in the cost of producing software in-house, without relying on creating dependencies or creating fewer dependencies. This can be modeled as an increase in the relative cost of a dependency. By simulating different magnitudes of the productivity improvement from AI-assisted coding, we show that the average systemic risk profile is flattened out by an increase of cost of merely five percent. The mechanism through which this risk reduction is achieved is the brute-force decrease in density for the network implied by increases in costs of dependencies. In general, our results help us to better understand the various driving forces and motifs in software development and how they shape the network of software dependencies.
@@
Our model is able to estimate the magnitude of externalities as well as homophily (Currarini et al., 2010; Jackson, 2008; Chandrasekhar, 2016), the tendency of individuals to form links to similar individuals. Furthermore, our model is able to detect heterophily (or competition) among maintainers. More specifically, we allow homophily to vary by unobservables, while in most models homophily is estimated only for observable characteristics (Currarini et al., 2010; Schweinberger and Handcock, 2015; DePaula et al., 2018; Chandrasekhar, 2016; Graham, 2020).
Lastly, our contagion analysis is broadly related to the body of work that treats security as a strategic game on a given graph. The literature following Goyal and Vigier (2014) and Dziubiński and Goyal (2013), for example, shows how the optimal point of attack of an adversary and the defense strategy of a designer are based on the centrality of the network. Differently from this literature, though, we also study how changing the structure of the network, induced by a changing relative cost of forming dependencies, affects the spread of contagion and find this to be highly effective.